Find the length of the arc on the unit circle with the given central angles 240º


1
Expert's answer
2021-01-20T13:19:05-0500

The length of the arc is

"s=r\theta"

where "r" is the radius and "\theta" is the central angle, in radians.

"240\degree=\frac{240}{180}\pi=\frac{4\pi}{3}" radians

The radius of the unit circle is 1.

"s=1\cdot\frac{4\pi}{3}=\frac{4\pi}{3}"

Answer: the length of the arc on the unit circle with the central angle of "240\degree" is "\frac{4\pi}{3}" units.